# HandS3C: 3D Hand Mesh Reconstruction with State Space Spatial Channel Attention from RGB images

### v3

@@ 4.3 Comparisons with the State-of-the-art Methods
We compare our proposed method with several recent state-of-the-art methods on three datasets as shown in Tables 1-3. As can be seen from Table 1, our method maintains state-of-the-art performance in the F@15 metric on a smaller number of datasets, although it is slightly worse than the other methods in the PA-MPJPE metric cho2022crossattention; 9710830. Moreover, its number of parameters shrinks by a factor of nearly 5 compared to models with higher accuracy cho2022crossattention. On the HO3D dataset, our method reduces the number of parameters by about 40% compared to recent methods Hampali2021KeypointTS. Meanwhile, when the number of parameters is similar to the method of Liu et al. liu2021semisupervised, all other metrics are greatly improved, and only the F-score metric is slightly reduced. To verify the effectiveness of our proposed method in the case of large-scale occlusion, the metrics on the DEXYCB dataset are shown in Table 2, and all the metrics outperform the recent state-of-the-art models Wang_2023; 10204388; Feng20243DHR. In conclusion, the experimental results on all three datasets show that our proposed method can still be effective in estimating hand postures when a large number of occlusions are included. Moreover, the state-of-the-art performance can be maintained while reducing the number of parameters, as shown in Fig.4 - 5.
@@
Figure 5: Qualitative results for DEXYCB. From left to right are the input image containing severe occlusions and angular distortions, the attention mapping map obtained by S3C, the heat map of keypoints in Regressor, the hand mesh and HandOCC park2022handoccnet hand mesh results.
Figure 6: Attention characterization map. It is worth noting that the S3C module is able to obtain sufficient hand information in the hand interaction edge region and the occlusion region. And the method expands the range of the receptive field so that it can capture the global feature mapping. (Orange is the occluded region and red is the hand interaction edge region)
4.4 Ablation Study
@@
In order to validate the effectiveness of the HandS3C method, we further analyze the different attention modules. The results of the ablation experiments on different datasets are shown in Tables 4-6. On the FreiHand and DEXYCB datasets, we perform three experiments, i.e., Baseline, Adding Self-Attention Module, and Adding S3C Module, as a way to validate that our proposed S3C is able to expand the effective receptive field while enhancing the hand features. On the HO3D dataset, we conducted five experiments: baseline, adding a self-attention module, adding a traditional state-space attention module (SSA), adding a serial state-space spatial-channel attention module (serial-S3C), and adding a parallel state-space spatial-channel attention module (S3C). With this approach, we demonstrate that a state-space model with added channel information enhances the ability to extract hand features under occlusion, with a wider effective receptive field, and the effectiveness of the parallel structure.
From Tables 4 and 5, we can see that in the presence of occlusion on the hand, the global feature information obtained from the self-attention mechanism, although it can enhance the hand features to a certain extent, will introduce more noise information. In Fig. 6, we visualize the attention distribution map, and it can be clearly seen that the feature map tends to characterize the global information after adding the attention mechanism, but still does not extract the hand features well for the occluded regions and hand interaction edges. The SSA module, on the other hand, aggregates more hand information and performs better at the interaction edges, but it will characterize the occluded part as the hand, which introduces interference information to a certain extent. The S3C module expands the range of the receptive field through the spatial plane selection mechanism, and then directionally strengthens the hand region features through the selection of the channel planes. Furthermore, by benefiting from linear directional scanning, the occluded areas are characterized using adjacent hand information, which reduces the introduction of interference through this global mapping capability.
In Table 6, it can be found that after the introduction of the state selection of the channel plane, the number of parameters of the model is only slightly increased by 0.01M, however, a significant improvement of nearly 2% is realized in the two performance metrics of PA-MPJPE and PA-MPVPE. This is due to the fact that the state selection in the channel plane enhances the selection and fusion of image features, and is able to extract more important channel information. The current state-space model focuses more on multi-directional scanning in the spatial plane, which is the reason for the 0.3% improvement in the F@5 metrics when using SSA. Using only the spatial plane for state selection, each pixel will focus more on neighboring pixels. However, observing other metrics will show that the channel dimension in the image will then contain richer feature information, which happens to be ignored by SSA.
@@
For the choice of serial and parallel structures, success was achieved in image restoration through serial structures in the study of Deng et al. Deng2024CUMambaSS. However, since SSM resembles the linear computation of recurrent neural networks (RNN). Serial scanning in space and channel add extra computation time. Meanwhile, compared to the parallel structure, the serial structure does not balance the contributions of spatial and channel dimensions well, as shown in the ablation experiments in Table 6.
5 Conclusion
@@
In order to effectively mitigate the challenges of the hand mesh reconstruction task under occlusion conditions, we introduced the state space model into the task of hand mesh reconstruction for the first time and propose the HandS3C method. The method expanded the effective receptive field and acquired hand features under occlusion by parallel cross-scanning in the spatial and channel planes. Among them, the spatial plane is able to acquire feature associations in different directions for each pixel, and the channel plane activates the target region features through state selection. This novel attention architecture is able to acquire sufficient global information and expand the effective sensory field while keeping the number of parameters small. Extensive experimental results show that this method is able to accurately predict hand postures from heavily occluded situations. Meanwhile, we expect this spatial and channel-wise state-space attention architecture to make progress in different downstream tasks.
[…]

### v4

@@ III-B Comparisons with the State-of-the-art Methods
We compare our proposed method with several recent state-of-the-art methods on three datasets as shown in Tables I-III. As can be seen in table I, our method only slightly underperforms the other methods[26] in the PA-MPJPE metrics, while remaining ahead in the other metrics, and it is important to note that the amount of parameters in our method has been reduced by a factor of almost 5[26]. On the HO3D dataset, our method reduces the number of parameters by about 40% compared to recent methods [4]. Meanwhile, when the number of parameters is similar to the method of Liu et al. [30], all other metrics are greatly improved, and only the F-score metric is slightly reduced. To verify the effectiveness of our proposed method in the case of large-scale occlusion, the metrics on the DEXYCB dataset are shown in Table II, and all the metrics outperform the recent state-of-the-art models [7, 9, 31, 32]. In conclusion, the experimental results on all three datasets show that our proposed method can still be effective in estimating hand postures when a large number of occlusions are included. Moreover, the state-of-the-art performance can be maintained while reducing the number of parameters, as shown in fig.2 - 3.
@@
Figure 3: Qualitative results for DEXYCB. From left to right are the input image containing severe occlusions and angular distortions, the attention mapping map obtained by S3C, the heat map of keypoints in Regressor, the hand mesh and HandOCC [5] hand mesh results.
III-C Ablation Study
@@
In order to verify the effectiveness of the HandS3C method, we further analyzed different attention modules. On the FreiHand and DEXYCB datasets, we conducted three experiments: baseline, adding a self-attention module, and adding an S3C module. On the HO3D dataset, we supplemented these three ablation experiments with experiments on the traditional state-space attention module (SSA) and the serial state-space spatial channel attention module (serial-S3C). With these ablation experiments, we demonstrate that the state-space model with added channel information can improve the ability to extract occluded partial hand features while extending the effective sensory field enhancement, and verify the effectiveness of the parallel structure. The results are summarized in Tables I-III.
From Tables I and III, it can be seen that on smaller scale datasets, the self-attention mechanism is difficult to learn the global features of the occluded part, but instead introduces more noise information to cause interference. The S3C module, on the other hand, expands the range of the sensory field through the spatial plane selection mechanism, and then enhances the hand region features directionally through the selection of the channel plane, which in turn avoids the interference of global noise. From Table III, it can be found that compared with the original SSM, the number of parameters of the model is only slightly increased by 0.01M after the introduction of the state selection of the channel plane, but a significant enhancement of nearly 2% is realized in the two performance metrics of PA-MPJPE and PA-MPVPE. This is due to the fact that our approach emphasizes the importance of channel information, which in combination with spatial information can expand the effective sensory field and thus extract the features of the region of focus.
@@
The study of Deng et al. successfully realized image restoration through serial structure [39]. However, since SSM resembles the linear computation of recurrent neural networks (RNN). Serial scanning in space and channel add extra computation time. Meanwhile, compared to the parallel structure, the serial structure does not balance the contributions of spatial and channel dimensions well, as shown in the ablation experiments in Table III.
IV Conclusion
@@
In order to effectively alleviate the challenges of the hand mesh reconstruction task under occlusion conditions, we firstly introduce the state space model into the hand mesh reconstruction task and propose the HandS3C method. The method enlarges the effective receptive field by performing parallel cross-scanning of the spatial plane and the channel plane. Thus, hand features under occlusion conditions are acquired while keeping fewer parameters. Extensive experimental results show that this approach can accurately predict hand pose under heavy occlusion. The ablation study also validates the effectiveness of the space state model in comparison to the attention-based model.
V Acknowledgment
This work was supported by the National Natural Science Foundation of the China, Shaanxi Provincial Key Industry Innovation Chain Program, 2024 public digital cultural service research project of the national public cultural development center of the Ministry of culture and tourism, and the scientific research program of Shaanxi Provincial Department of Education [grant numbers 62072362 and 12101479, 2020ZDLGY07-05, GGSZWHFW2024-017, 23jp060].